Which expression can be substituted for x in the top equation of the following system to solve the system by substitution? What is the solution?
−2x−3y=−13
x+y=6
The choices are :
1) −y+6, (5,1)
2) y+6, (-1/5 ,1/5) that's meaning 1 over neg 5,and 1 over 5

3)-3/2y+13/2 (5,1)
4) 6−y, (1,5)

Answer by Theo(13342)

the equations to solve are:

−2x − 3y = −13

x + y = 6

in the second equation, solve for x to get x = -y + 6.

in the first equation, replace x with (-y + 6) to get -2(-y+6) -3y = -13.

simplify to get 2y - 12 - 3y = -13

combine like terms to get -y - 12 = -13

solve for y to get y = 1.

go back to your original equations.

they are:

−2x − 3y = −13

x + y = 6


replace y in the second equation with 1 to get x + 1 = 6

solve for x to get x = 5.

your solution is x = 5 and y = 1.

replace x with 5 and y with 1 in the first original equation to get −2(5) − 3(1) = −13

simplify to get -10 - 3 = -13 which reduces to -13 = -13 which is true.

this conforms the solution is correct.

your solution is (x,y) = (5,1).

your choices are:

1) −y+6, (5,1)
2) y+6, (-1/5 ,1/5) that's meaning 1 over neg 5,and 1 over 5
3)-3/2y+13/2 (5,1)
4) 6−y, (1,5)

eliminate selection 2 and 4 because their coordinate solution is not (5,1).

that leaves selection 1 or selection 3.

you were asked to substitute for x in the first equation.

you did that by solving for x in the second equation and then substituting for x in the first equation.

when you solved for x in the second equation, you got x = -y + 6.

you solution is therefore selection 1.

selection 1 is:

1) −y+6, (5,1)

the first part of that selection is the value that you substituted for x.

the second part of that selection is the solution in coordinate form.